# Single color virtual H&E staining with In-and-Out Net

### v1

@@ Abstract
Virtual staining streamlines traditional staining procedures by digitally generating stained images from unstained or differently stained images. While conventional staining methods involve time-consuming chemical processes, virtual staining offers an efficient and low infrastructure alternative. Leveraging microscopy-based techniques, such as confocal microscopy, researchers can expedite tissue analysis without the need for physical sectioning. However, interpreting grayscale or pseudo-color microscopic images remains a challenge for pathologists and surgeons accustomed to traditional histologically stained images. To fill this gap, various studies explore digitally simulating staining to mimic targeted histological stains. This paper introduces a novel network, In-and-Out Net, specifically designed for virtual staining tasks. Based on Generative Adversarial Networks (GAN), our model efficiently transforms Reflectance Confocal Microscopy (RCM) images into Hematoxylin and Eosin (H&E) stained images. We enhance nuclei contrast in RCM images using aluminum chloride preprocessing for skin tissues. Training the model with virtual H&E labels featuring two fluorescence channels eliminates the need for image registration and provides pixel-level ground truth. Our contributions include proposing an optimal training strategy, conducting a comparative analysis demonstrating state-of-the-art performance, validating the model through an ablation study, and collecting perfectly matched input and ground truth images without registration. In-and-Out Net showcases promising results, offering a valuable tool for virtual staining tasks and advancing the field of histological image analysis.
@@
virtual staining , reflectance confocal microscopy (RCM) , Generative adversarial network (GAN) , In-and-Out Net
@@ 1 Introduction
Virtual staining refers to methods that use computational models to generate histological stained images from other images. In histology image analysis, the traditional approach is to use dyes to chemically stain frozen sectioned or formalin fixed paraffin embedded tissue block sections like Hematoxylin and Eosin (H&E) staining. This fixing-sectioning-staining-diagnosing process is time-consuming and high infrastructure demanding. For instance, in some skin surgeries such as Mohs microscopic surgery, cancerous tissue is removed in an iterative process, with each layer removal referred to as a stage. For Mohs surgery, the whole process from fixation to diagnosis can require at least 30-60 mins for each stage, and requires skilled technicians and substantial equipment. To enhance efficiency and simplify these kinds of procedures, microscope-based methodsLongo et al. (2014); Patel et al. (2007); Abeytunge et al. (2013); Yoshitake et al. (2018); Glaser et al. (2017); A. Austin et al. (2016); Cui et al. (2018) offer a practical substitute for traditional histological staining. By swiftly staining fresh tissueChen et al. (2021) or even eliminating the staining step altogetherRivenson et al. (2019), these approaches bypass certain time-consuming and infrastructure-intensive stages. Moreover, they enable imaging of bulk tissue without the need for physical sectioningBoktor et al. (2022). Consequently, these methods not only expedite the process but also require substantially less infrastructure, usually limited to a single microscope deviceBoktor et al. (2022).
@@
Despite the efficiency of the microscopic approaches, pathologists and surgeons are accustomed to interpreting histologically stained images rather than grayscale or pseudo-color microscopic images. Addressing this challenge, numerous studiesBai et al. (2023) have explored using computers to digitally simulate the staining of microscopic images to resemble targeted histological stained images. Furthermore, there is a growing demand for obtaining various types of stained images from a single tissue sample. For instance, researchers may require both H&E and Masson’s Trichrome (TRI) stained images from the same live biopsy section to examine detailed structures comprehensivelyFoot (1933). This necessity is met by stain-to-stain virtual staining, a technique that leverages computers to transform images from one histological stain domain to anotherBai et al. (2023); de Haan et al. (2021). The process of converting unstained microscopic grayscale or pseudo-color images to histologically stained images is referred to as ”label-free” virtual staining. In contrast, labeled virtual staining involves staining samples with chemicals to convert them into pseudo-color histological images. Additionally, transforming images from one histological stain domain to another is termed ”stain-to-stain” virtual staining. While these techniques serve different purposes, they share similar underlying approaches.
@@
The initial exploration into virtual staining introduced a method that involves the linear combination of two fluorescence confocal microscope channels to produce H&E-like images.Bini_H&E; Gareau et al. (2009) One channel of a microscope only collects lights within a fixed range of wavelength. So, one channel always represents one color in a microscopic images. In this technique, researchers employed two distinct dyes for staining nuclei and collagen/cytoplasm, with each dye corresponding to a separate channel. Different R, G, B weights were assigned to each channel, enabling their amalgamation into an RGB image. Subsequently, Beer’s law was applied to nonlinearly combine these channels, enhancing the realism of the resulting H&E image. It is noteworthy that both channels need not originate from the fluorescence modality exclusively. Alternatively, a reflectance channel can be used to depict collagen/cytoplasm, while another fluorescence channel highlights nuclei.Gareau (2009); Daniel Gareau et al. (2012) As long as one channel captures collagen/cytoplasm and the other showcases nuclei, this methodology remains applicable. This versatile approach is sometimes employed to establish ground truths for virtual staining across various modalities. For instance, when virtually staining reflectance confocal images (RCM) to H&E images, Li et. alLi et al. (2021) generated H&E ground truths from two channels, utilizing the approach to circumvent the need for actual chemical H&E staining.
@@
The rise of artificial intelligence has spurred the development and deployment of numerous generative deep learning models in the innovative field of virtual staining. Among the array of generative models, the prevalent choice for virtual staining applications is the utilization of Generative Adversarial Networks (GAN)Goodfellow et al. (2014). Ozcan’s laboratory, in particular, has spearheaded groundbreaking research in this domain. Riverson et al.Rivenson et al. (2019, 2019) used pix2pixGANIsola et al. (2017) to conduct some remarkable and clinically significant label-free virtual staining studies. They virtually stain Quantitative Phase Images (QPI)Rivenson et al. (2019) and auto-fluorescence imagesRivenson et al. (2019) into multiple histological stained counterparts, including H&E, TRI stain and Jones methenamine silver (JMS) stain. The tissue types they used included human skin, gland, kidney, liver, thyroid, and lung. Continuing Riverson’s work, De Hann et al.de Haan et al. (2021) pioneered the use of CycleGANZhu et al. (2017) in stain-to-stain tasks, enabling the transformation of H&E-stained images into a diverse array of histological stains, including TRI stain, JMS stain and Periodic acid–Schiff (PAS) stain. Furthermore, Yang et al.Yang et al. (2022) combined the techniques of Riverson and De Hann to employ Cascaded GAN networks to virtually stain auto-fluorescence images into PAS-stained images. They first conducted label-free staining using pix2pixGAN to stain auto-fluorescence images into H&E-stained images, then they used CycleGan to stain H&E images into PAS-stained images. They proved this two-step approach achieved better performance than directly stained auto-fluorescence images into PAS-stained images. The works done by Riverson, De Hann and Yang et al. excellently exemplify the outstanding applications of GAN-based models in the virtual staining field. These models are capable of virtually staining many diverse human tissue samples and can generate various distinct histological stain styles. Beyond the realm of GAN-based models, noteworthy strides have been made in other generative models. Lotfollahi et al.Lotfollahi et al. (2019) employed a straightforward encoder-decoder architecture, transforming Fourier transform infrared (FTIR) spectroscopic images into H&E and DAPI (4,6-diamidino-2-phenylindole) images. Burlingame et al.Burlingame et al. (2020) used a Variational Autoencoder (VAE)Kingma and Welling (2022) network to stain H&E-stained images into immunofluorescent images. Researchers also made efforts on modifying state-of-the-art models like GAN-based models to get better performance on virtual staining tasks. Li et al.Li et al. (2021) integrated attention blocks into the pix2pixGAN model, specifically tailoring it to stain RCM images into the H&E-stained images. This marked a successful attempt to modify the network for the virtual staining tasks. Additionally, Liu et al.Liu et al. (2023) introduced the VSGD-Net for staining H&E to Sox10 images, exhibiting outstanding performance in melanocyte detection. These diverse approaches collectively contribute to the multifaceted landscape of generative models, expanding the possibilities in virtual staining research.
@@
While state-of-the-art models have garnered notable success in virtual staining, there has been a limited number of models explicitly designed for virtual staining tasks. This paper addresses this gap by introducing a novel network crafted for the purpose of virtual staining. The effectiveness of our proposed model is demonstrated through its successful application in staining RCM images to produce H&E-stained images.
@@
This study focuses on employing GAN-based models to achieve grayscale-to-H&E virtual staining. Over the past few decades, diverse microscope-based techniques, including confocal microscopyLongo et al. (2014); Patel et al. (2007); Abeytunge et al. (2013), MUSEYoshitake et al. (2018), light-sheet microscopyGlaser et al. (2017), CARSA. Austin et al. (2016); Cui et al. (2018), and SRS microscopyA. Austin et al. (2016); Cui et al. (2018), have been investigated to speed up the Mohs tissue processing. Among these, RCM stands out as a leading optical imaging method with a commercial device (VivoScope, Caliber I.D.) on the market with US reimbursement codes, making it a forefront contender for surgical guidance development. However, the utilization of RCM for virtual staining has been minimally explored due to its limited bright nuclei contrast and its single-channel nature. RCM captures information in a unified manner, making it difficult to separate nuclei and collagen/cytoplasm images. A significant contribution in this domain comes from the work of Li et al.Li et al. (2021) They pioneered a two-stage digital staining process for RCM images. Initially, they transformed original RCM images into acetic acid-stained RCM images, enhancing the bright nuclei contrast. Subsequently, they extended this methodology by digitally staining acetic acid-stained RCM images to simulate H&E images, presenting an innovative approach within the field.
@@
This study introduces a GAN-based architecture to transform single channel RCM images into H&E-stained images. To enhance nuclei contrast under RCM, we employed aluminum chloride as a preprocessing agent for skin tissues. To train the model, we utilized virtual H&E labels with two fluorescence channels to provide pixel level ground truth without needing to perform image registration. Our contributions can be succinctly summarized as follows:
@@
    Introduction of In-and-Out Net: We proposed the In-and-Out Net, a model meticulously designed for the type of label-free virtual staining tasks that the training ground truth H&E images were synthesized through two channels. It is also applicable when the ground truth images are chemically stained and can be decomposed into two channels representing different dyes. The model is structured with an inner loop and an outer loop, each serving distinct roles in the virtual staining process. Additionally, we propose an optimal training strategy for the model.
@@
    Comparative Analysis: We conducted a comprehensive comparison between our model and previous approaches for both label-free and stain-to-stain virtual staining. The results demonstrate that our model achieves state-of-the-art performance.
@@
    Model Validation: We substantiate the essentialness and appropriateness of our model architecture through an ablation study, reinforcing its effectiveness.
@@
    Data Collection: We collected reflectance channel images and two channels of fluorescence images using the same microscope, creating a perfect pixel-to-pixel match between input and ground truth images without the need for registration.
@@ 2 Methods and experiments
In this section, we describe our data collection, data preprocessing, the design of our proposed In-and-Out Net, and the training approach. In the ”Results and discussion” section, we compare our model results with several state-of-the-art models and ablate In-and-Out Net’s components to validate its design. We also compare the generator loss within different training approaches.

### v2

# Single color digital H&E staining with In-and-Out Net

@@ Abstract
Digital staining streamlines traditional staining procedures by digitally generating stained images from unstained or differently stained images. While conventional staining methods involve time-consuming chemical processes, digital staining offers an efficient and low-infrastructure alternative. Researchers can expedite tissue analysis without physical sectioning by leveraging microscopy-based techniques, such as confocal microscopy. However, interpreting grayscale or pseudo-color microscopic images remains challenging for pathologists and surgeons accustomed to traditional histologically stained images. To fill this gap, various studies explore digitally simulating staining to mimic targeted histological stains. This paper introduces a novel network, In-and-Out Net, designed explicitly for digital staining tasks. Based on Generative Adversarial Networks (GAN), our model efficiently transforms Reflectance Confocal Microscopy (RCM) images into Hematoxylin and Eosin (H&E) stained images. Using aluminum chloride preprocessing for skin tissue, we enhance nuclei contrast in RCM images. We trained the model with digital H&E labels featuring two fluorescence channels, eliminating the need for image registration and providing pixel-level ground truth. Our contributions include proposing an optimal training strategy, conducting a comparative analysis demonstrating state-of-the-art performance, validating the model through an ablation study, and collecting perfectly matched input and ground truth images without registration. In-and-Out Net showcases promising results, offering a valuable tool for digital staining tasks and advancing the field of histological image analysis.
@@
digital staining , reflectance confocal microscopy (RCM) , Generative adversarial network (GAN) , In-and-Out Net
@@ 1 Introduction
Digital staining refers to methods that use computational models to generate histological stained images from other images. In histology image analysis, the traditional approach uses dyes to chemically stain frozen sectioned or formalin fixed paraffin embedded tissue block sections like Hematoxylin and Eosin (H&E) staining. In clinical practice, histological stains such as Hematoxylin and Eosin (H&E) are vital for the microscopic examination of tissue samples, allowing pathologists to diagnose diseases like cancer, identify infection, and assess tissue abnormalities. However, this fixing-sectioning-staining-diagnosing process is time-consuming and highly infrastructure-demanding. For instance, in some skin surgeries such as Mohs microscopic surgery, cancerous tissue is removed in an iterative process, with each layer removal referred to as a stage. For Mohs surgery, the process from fixation to diagnosis can require at least 30-60 mins for each stage, requiring skilled technicians and substantial equipment.
To eliminate the need for fixation and staining, microscope-based methods(Longo et al. (2014); Patel et al. (2007); Abeytunge et al. (2013); Yoshitake et al. (2018); Glaser et al. (2017); Austin et al. (2016); Cui et al. (2018)) offer a practical substitute for traditional histological staining. By swiftly staining fresh tissue(Chen et al. (2021)) or even eliminating the staining step altogether(Rivenson et al. (2019a)), these approaches bypass certain time-consuming and infrastructure-intensive stages. Moreover, they enable bulk tissue imaging without the need for physical sectioning(Boktor et al. (2022)). Consequently, these methods expedite the process and require substantially less infrastructure, usually requiring only a single microscope device(Boktor et al. (2022)).
@@
Despite the efficiency of the microscopic approaches, they have not been widely adopted into clinical practice. This is likely because their visual output looks so different than traditional pathology, that the clinicians can’t utilize them without additional training. Pathologists and surgeons are accustomed to interpreting colored histologically- stained images rather than grayscale or pseudo-color microscopic images. Addressing this challenge, numerous studies(Bai et al. (2023)) have explored using computers to digitally simulate the staining of microscopic images to resemble targeted histological stained images. Digital staining offers a practical alternative by allowing for the generation of stained-like images without chemical stains. Furthermore, there is a growing demand for obtaining various types of stained images from a single tissue sample. For instance, researchers may require both H&E and Masson’s Trichrome (TRI) stained images from the same live biopsy section to examine detailed structures comprehensively(Foot (1933)). This necessity was met by stain-to-stain digital staining, a technique that leverages computers to transform images from one histological stain domain to another(Bai et al. (2023); de Haan et al. (2021)). The process of converting unstained microscopic grayscale or pseudo-color images to histologically stained images is referred to as ”label-free” digital staining. In contrast, labeled digital staining involves converting samples with chemicals into pseudo-color histological images. Additionally, transforming images from one histological stain domain to another is termed ”stain-to-stain” digital staining. While these techniques serve different purposes, they share similar underlying approaches.
In conclusion, digital staining aims to produce images that resemble traditional histologically stained slides in a fraction of the time. Techniques like RCM can be used in real-time surgical settings, where quick decision-making is crucial. Digital staining eliminates several steps in the traditional process, reducing time and infrastructure requirements. In these scenarios, a microscope and computational model can replace the complex staining procedures, making digital staining an attractive option for use in operating rooms or point-of-care diagnostics.
Another significant advantage is the ability of digital staining to address non-invasive diagnostic needs. Pathologists and clinicians are accustomed to interpreting histologically stained images, which provide a wealth of diagnostic information. However, in procedures where tissue biopsies are impractical, such as in vivo imaging using techniques like RCM or optical coherence tomography (OCT), the grayscale or pseudo-color images produced by these modalities are often challenging to interpret. Digital staining bridges this gap by converting these grayscale images into colorized images that resemble traditional H&E-stained tissue. This enhances the diagnostic utility of non-invasive imaging, especially for identifying conditions like basal cell carcinoma or melanoma without the need for a physical biopsy.
Digital staining also has the potential to improve telepathology and remote diagnostics. In resource-limited regions, removed skin tissue must be packaged in formalin, shipped to a lab, then processed and stained. Sometimes, this histopathology would then need to be scanned into a digital image and sent to a pathologist at another site. By acquiring digital images in the clinic that can be quickly sent to anyone in the world for interpretation, several steps (and costs) can be eliminated, making diagnostic workflows faster and more accessible. This is especially valuable for remote consultations where physical slide preparation and transport would cause delays.
@@
The initial exploration into digital staining introduced a method that involves the linear combination of two fluorescence confocal microscope channels to produce H&E-like images.(Bini et al. (2011); Gareau et al. (2009)) Specifically, one channel mimics the hematoxylin (H) stain, while the other mimics the eosin (E) stain, with each channel capturing light within a specific wavelength range to represent a distinct color in the final image. In this technique, researchers employed two distinct dyes for staining nuclei and collagen/cytoplasm, each corresponding to a separate channel. Each channel was assigned different R, G, and B weights, enabling their amalgamation into an RGB image. Subsequently, Beer’s law was applied to nonlinearly combine these channels, enhancing the realism of the resulting H&E image. Notably, channels can be independent of the fluorescence modality. Alternatively, a reflectance channel can be used to depict collagen/cytoplasm, while another fluorescence channel highlights nuclei.(Gareau (2009); Daniel Gareau et al. (2012)) As long as one channel captures collagen/cytoplasm and the other showcases nuclei, this methodology remains applicable. This versatile approach sometimes establishes ground truths for digital staining across various modalities. For instance, when digitally staining RCM to H&E images, Li et al. (Li et al. (2021)) generated H&E ground truths from two channels, utilizing the approach to circumvent the need for actual chemical H&E staining.
@@
The rise of artificial intelligence has spurred the development and deployment of numerous generative deep-learning models in the innovative field of digital staining. Among the array of generative models, the prevalent choice for digital staining applications is the utilization of Generative Adversarial Networks (GAN)(Goodfellow et al. (2014)). Ozcan’s laboratory, in particular, has spearheaded groundbreaking research in this domain. Riverson et al.(Rivenson et al. (2019a, b)) used pix2pixGAN(Isola et al. (2017)) to conduct some remarkable and clinically significant label-free digital staining studies. They digitally stain Quantitative Phase Images (QPI)(Rivenson et al. (2019a)) and auto-fluorescence images(Rivenson et al. (2019b)) into multiple histological stained counterparts, including H&E, TRI stain and Jones methenamine silver (JMS) stain. Their tissue types included human skin, gland, kidney, liver, thyroid, and lung. Continuing Riverson’s work, De Hann et al.(de Haan et al. (2021)) pioneered the use of CycleGAN(Zhu et al. (2017)) in stain-to-stain tasks, enabling the transformation of H&E-stained images into a diverse array of histological stains, including TRI stain, JMS stain and Periodic acid–Schiff (PAS) stain. Furthermore, Yang et al.(Yang et al. (2022)) combined the techniques of Riverson and De Hann to employ Cascaded GAN networks to stain auto-fluorescence images into PAS-stained images digitally. They first conducted label-free staining using pix2pixGAN to stain auto-fluorescence images into H&E-stained images; then, they used CycleGan to stain H&E images into PAS-stained images. They proved this two-step approach performed better than directly stained auto-fluorescence images into PAS-stained images. The works done by Riverson, De Hann, and Yang et al. excellently exemplify the outstanding applications of GAN-based models in the digital staining field. These models can digitally stain diverse human tissue samples and generate various histological stain styles. Beyond GAN-based models, noteworthy strides have been made in other generative models. Lotfollahi et al.(Lotfollahi et al. (2019)) employed a straightforward encoder-decoder architecture, transforming Fourier transform infrared (FTIR) spectroscopic images into H&E and DAPI (4,6-diamidino-2-phenylindole) images. Burlingame et al.(Burlingame et al. (2020)) used a Variational Autoencoder (VAE)(Kingma and Welling (2022)) network to stain H&E-stained images into immunofluorescent images. Researchers also tried modifying state-of-the-art models, like GAN-based models, to perform better on digital staining tasks. Li et al.(Li et al. (2021)) integrated attention blocks into the pix2pixGAN model, tailoring them to stain RCM images into the H&E-stained images. This marked a successful attempt to modify the network for the digital staining tasks. Additionally, Liu et al.(Liu et al. (2023)) introduced the VSGD-Net for staining H&E to Sox10 images, exhibiting outstanding performance in melanocyte detection. These diverse approaches collectively contribute to the multifaceted landscape of generative models, expanding the possibilities in digital staining research.
@@
While state-of-the-art models have garnered notable success in digital staining, only a limited number of models are explicitly designed for digital staining tasks. This paper addresses this gap by introducing a novel network crafted for digital staining. The effectiveness of our proposed model is demonstrated through its successful application in staining RCM images to produce H&E-Stained images.
@@
This study focuses on employing GAN-based models to achieve grayscale-to-H&E digital staining. Over the past few decades, diverse microscope-based techniques, including confocal microscopy(Longo et al. (2014); Patel et al. (2007); Abeytunge et al. (2013)), MUSE(Yoshitake et al. (2018)), light-sheet microscopy(Glaser et al. (2017)), CARS(Austin et al. (2016); Cui et al. (2018)), and SRS microscopy(Austin et al. (2016); Cui et al. (2018)), have been investigated to speed up the Mohs tissue processing. Among these, RCM stands out as a leading optical imaging method with a commercial device (VivoScope, Caliber I.D.) on the market with US reimbursement codes, making it a forefront contender for surgical guidance development. However, the utilization of RCM for digital staining has been minimally explored due to its limited bright nuclei contrast and its single-channel nature. RCM captures information in a unified manner, making it challenging to separate nuclei and collagen/cytoplasm images. A significant contribution in this domain comes from the work of Li et al.(Li et al. (2021)). They pioneered a two-stage digital staining process for RCM images. Initially, they transformed original RCM images into acetic acid-stained RCM images, enhancing the bright nuclei contrast. Subsequently, they extended this methodology by digitally staining acetic acid-stained RCM images to simulate H&E images, presenting an innovative approach within the field.
@@
This study introduces a GAN-based architecture to transform single-channel RCM images into H&E-stained images. To enhance nuclei contrast under RCM, we employed aluminum chloride as a preprocessing agent for skin tissues. To train the model, we utilized digital H&E labels with two fluorescence channels to provide pixel-level ground truth without needing to perform image registration. Our contributions can be succinctly summarized as follows:
@@
    Introduction of In-and-Out Net: We proposed the In-and-Out Net, a model meticulously designed for label-free digital staining tasks where the training ground truth H&E images are synthesized through two channels. It also applies when the ground truth images are chemically stained and can be decomposed into two channels representing different dyes. The model is structured with an inner loop and an outer loop, each serving distinct roles in the digital staining process. Specifically:
    • (a)
      Inner Loop (Detailed Morphological Big-Tuning): The inner loop is responsible for the detailed morphological big-tuning of the image. This loop focuses on accurately capturing large-scale morphological features, such as cell boundaries, nuclei, and overall tissue structure. It operates primarily on the structural aspects of the image, ensuring that the major morphological elements are correctly represented before moving on to finer adjustments. By splitting the image into separate H and E channels, the inner loop allows the model to isolate and refine important structures better, ensuring that these morphological details are clear and precise.
    • (b)
      Outer Loop (Fine-Tuning of Morphology and Color): The outer loop performs fine-tuning, not only on the morphological details but also on the color properties of the image. Once the inner loop has established a strong foundation of structural information, the outer loop adjusts the finer details of the morphology, such as smoothing out boundaries, refining textures, and ensuring small-scale accuracy. In addition, the outer loop handles the accurate colorization of the image, combining the H and E channels to create a final output with the realistic colors and contrast expected from H&E staining. This step is crucial in ensuring that the color harmonizes well with the morphology, as accurate color is essential for diagnostic purposes in histology.
@@
    Tailored for digital H&E Staining: Unlike traditional image-to-image translation models, our model is specifically designed to address the unique challenges of digital staining. The in-and-out loop framework ensures that morphological and color properties are learned separately but integrated smoothly for better accuracy in the final output.
@@
    Comprehensive Comparative Analysis and Ablation Study: We comprehensively compared our model and previous approaches for label-free and stain-to-stain digital staining. The results demonstrate that our model achieves state-of-the-art performance. Additionally, we performed an extensive ablation study to validate the contribution of each component in the architecture. The study shows that combining the inner loop, outer loop, and H&E channel separation leads to significantly improved image quality, both in structure and color.
@@
    Data Collection and Model Validation: We collected reflectance channel images and two channels of fluorescence images using the same microscope, creating a pixel-to-pixel match between input and ground truth images without the need for registration. This precise data collection, combined with the tailored architecture of our model, ensures the production of highly accurate digital staining results.
  5. 5.
    Optimized for Hierarchical Learning: The model is optimized to handle both the structure-color interaction and the complexity of histological images, ensuring that even complex tissue morphologies are accurately stained digitally. This method addresses the specific needs of pathology and histological image analysis, maintaining high structural fidelity and realistic color mapping.
@@ 2 Methods and experiments
In this section, we describe our data collection, data preprocessing, the design of our proposed In-and-Out Net, and the training approach. In the Results and discussion section, we compare our model results with several state-of-the-art models and ablate In-and-Out Net’s components to validate its design. We also compare the generator loss within different training approaches.